David D. Golombek

Term Address Home Address 3 Ames St, Box 49 1078 E. Balmoral Place Cambridge, MA 02142 Tuscon, AZ 85737 (617) 225-6532 (520) 825-2121 Email: daveg@mit.edu

Objective A position where I can gain more experience in the design of computer systems and continue to learn more about the construction of individual components.

Education Massachusetts Institute of Technology . . . . . . . . . . . . . . . . . . . . . . . . . . . 1994—present Pursuing a Bachelor’s Degree in Computer Science and Engineering with a minor in Philosophy. Completed the following courses: Structure and Interpretation of Computer Programs (6.001), Circuits and Electronics (6.002), other non-major related courses.

William S. Hart High School . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1990—1994 Valedictorian, 4.53 Cumulative GPA.

Extracurricular activities at MIT include membership on the Student Information Processing Board (SIPB), a student group that encourages the use of computers by students on campus, and membership on the men’s gymnastics team.

Experience Integrated Computer Engines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Waltham, MA Software Engineer October 1995—present Wrote multi-threaded, multi-process simulator of parallel computer, and ported from OS/2 to UNIX and Windows NT and 95. Also wrote a data visualization tool using MFC, Eisa and PCI device drivers for Windows NT, a SHARC assembly language diagnostic utility for low level hardware verification, a windowed front end to all tools offered by company, and designed a Porting API for flexibility in moving all company tools between operating systems.

Cambridge Technology Group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Cambridge, MA Programmer, Internet Specialist January 1995—October 1995 Helped develop WWW home page, including CGI script development, page creation, and layout design. Joined other projects as needed, including Perl programming, recruitment and hiring, and project demonstrations.

Jet Propulsion Laboratory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Pasadena, CA Technical Hire June 1993—September 1994 Research and study of Planetary Geology, performing digital image processing, manipulation, and measuring, on MicroVax and Sun Workstations. Worked as assistant system manager on MicroVax and wrote/modified programs as needed.

Independent Computer Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Valencia, CA Computer Consultant June 1993—present Private work for individuals, repairing, installing, and maintaining personal computers. Custom programming of Windows Database under Paradox 4.5, shareware programming with C, Turbo Pascal, and Turbo Assembler.

Taos Ski Valley, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .Taos, NM General Hire December-January 1991, 1992, 1993 Temporary hire at ski valley for winter rush, assisting in impacted areas.

Six Flags Magic Mountain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .Valencia, CA Ride Operations June 1992—November 1992 Operated rollercoasters at local amusement park.

Appointed USA Today All-American Academic Third Team, June 1994.

Honors and Awards Awarded the National Council of Teachers of English Excellence in Writing Award, 1994. Self-taught knowledge of Unix, VMS, DOS, Windows (NT, 95, 3.1), and Macintosh machines. Self-taught programming in C, C++, Java, Pascal, Lisp, TCL/tK, perl, and Assembly for the x86 and SHARC. Self-taught knowledge of the MFC, Win32 API, HTML, and LaTeX. Eagle Scout.

References Available upon request.